Factor the following perfect square trinomial.
64z^2+16z+1
I always have problems with me, can someone please explain. Thank you!

Found 2 solutions by Alan3354, Chelzhampton :
Answer by Alan3354(69443)

Factor the following perfect square trinomial.
64z^2+16z+1
-----------------
If you know it's a perfect square, it's the sqrt of the end terms.

Answer by Chelzhampton (1)

Knowing that 8 * 8 = 64, and
8+ 8= 16,
(8z + 1)(8z + 1)
To explain,
Foil- and then you get
64z^2+8z+8z+1
The 8z times 8z makes 64z^2
8z & 8z add making 16z
You're left with 64z^2 +16z+1
Hopefully that explains it